# An AI System for Continuous Knee Osteoarthritis Severity Grading Using Self-Supervised Anomaly Detection with Limited Data

### v1

@@ 5 Implementation Details
The hyper-parameters including the model backbones were selected based on the model’s performance on the validation set. The models were trained with a batch size of one, a learning rate of $061e-$ 06, Adam optimiser [21] and weight decay of $0.1$.
5.1 Stage 1
The size of the training set for each ensemble, was set to 30 ($N=30$) as per the original FewSOME [7] implementation. The size of the training data, $X_{train}$ (the data where $N$ is sampled from for each ensemble) was set to 150 and a total of $K=10$ ensembles were trained. Section 7.3 presents an analysis to motivate the selection of a training set size of 150 examples. The first five layers of an AlexNET [22] was employed as the model backbone. Training was halted once the training loss began to plateau.
5.2 Stage 3iter1
As X-rays with psuedo labels are now available for training in stage 3, the larger VGG-16 [33] is employed. The margin, $m$ was set to $1.184$ to obtain the pseudo labels for DCRL-FSOA as this resulted in a balanced number of normal instances and anomalous instances (30 pseudo labels). Based on experiments on validation data, the margin $m$ was set to $3.122$ to obtain the pseudo labels to train DCRL-FSsev. The threshold $t$ for combining the models was set to the AS of the 95% percentile of the training data as output by DCRL-FSsev. The training is halted when the CD between the normal centre, $C_{norm}$ and anomalous centres $C_{OA}$ (when training DCRL-FSOA) or $C_{sev}$ (when training DCRL-FSsev) begins to plateau. Section 7.4 presents an analysis of the effects of this early stopping technique.
5.3 Stage 3iterfinal
As previously mentioned, DCRL-FSOA can be trained iteratively. In the second iteration, the number of pseudo labels was equal to $263$ for DCRLOA (with $m=1.184$ as before). To avoid a large class imbalance, the number of training instances for each ensemble was increased from $30$ to $150$ (all available labelled training data). The model was not trained for further iterations as it reached optimal performance.
@@ 6 Results
The methodology is assessed based on its ability to diagnose OA according to the KL scale (grades $\geq 2$) and the OARSI scale. Table 2 reports the Area Under the Receiver Operator Curve (AUC) in % for both tasks, denoted as $AUC_{KL}$ and $AUC_{O}$. The Spearman Rank Correlation Coefficient (SRC) between the model output and the KL grade is also reported as $SRC_{KL}$. Finally, to ensure that the model can identify severe OA cases to a high degree of accuracy, the table reports the performance of detecting cases with KL grades $>3$ as $AUC_{KL_{g>3}}$.
@@ 6.1 Stage 1
Each method had $150$ training instances available for training. The proposed SS-FewSOME significantly outperforms competing methods and it shows substantial performance improvement on the original FewSOME implementation, increasing the $AUC_{KL}$ from $58.3$% to $72.9$%. The results also show that integrating the patches method into the model has resulted in a performance improvement, increasing $SRC_{KL}$ from 0.315 to 0.432.
@@ 6.2 Stage 3iter1
The trade-off between OA detection and severe OA detection accuracy can be evidently seen from the results of DCRL-FSOA (stage $3_{iter1}$) and DCRL-FSsev, with the former outperforming DCRL-FSsev in terms of OA detection but underperforming in terms of severe OA detection. However, the combined model DCRL-FScomb (stage $3_{iter1}$) combines both aspects of the models to accurately detect OA, severe OA and correlate with the KL grades with an SRC of 0.541. Figure 5 shows the distance of each data instance in the test set from $C_{norm}$ at stage 1 and stage 3iter1. Separation between classes can be evidently seen at stage 1 and furthermore at stage 3iter1. Figure 6 presents the most anomalous examples as output by DCRL-FScomb (stage $3_{iter1}$). The common symptoms of OA are apparent such as JSN and the presense of osteophytes.
@@ 6.3 Stage 3iterfinal
The final iteration achieved an SRC of 0.583, a promising result given the inter-rater reliability of human experts was reported to be as low as 0.51 in some cases [37]. The proposed methodology outperforms the competing method DDAD [11] by margins of up to 24%, which was trained on $150$ labelled healthy knee X-rays and 5,628 unlabelled X-rays. Notably, the proposed work outperforms the previous OA automated continuous system that reports an SRC of 0.524 (averaged across classes), having trained on 10,012 labelled images [24]. Additionally, a paired student t-test showed that the difference between the mean model output for each KL grade was statistically significant at the 5% significance level.
@@ 7 Further Analysis
7.1 CLIP Ablation
@@
DCRL-FScomb 3iter1 $\mathbf{78.6\pm{}1.4}$ $\mathbf{79.5\pm{}1.2}$ $\mathbf{97.2\pm{}0.4}$ $\mathbf{0.541\pm{}0.0}$
@@ 7.4 Analysis of Early Stopping
8 Conclusion
This work has demonstrated the potential for a continuous disease severity grading system, which holds promise for substantial clinical impact by eliminating subjectivity and enabling the assessment of disease severity at any stage along a continuum, rather than predefined discrete categories. This method is particularly useful for capturing subtle variations and changes over time, providing a more detailed and dynamic understanding of the condition’s progression, ultimately resulting in improved patient care. The analysis showed that the proposed method outperforms existing methods by significant margins and can achieve human level performance whilst training on few labels, thus eliminating the requirement for large annotated databases.

### v2

@@ 5 Implementation Details
The implementation details are summarised in table 2.
Hyper-Parameters. The hyper-parameters and model backbones that had the best performance on the validation set were selected. The models were trained with $K=10$, learning rate of $061e-$ 06, batch size of one, weight decay of $0.1$ and the Adam optimiser [25]. The first five layers of the AlexNET [26] was used as the model backbone for Stage 1, while the larger VGG-16 [40] was implemented for later stages where more training data was available. As previously mentioned, DCRL-FSOA can be trained iteratively, meaning it can be used to predict additional pseudo labels, which are denoised using the method of Stage 2 and then retrained on the newly assigned pseudo labels. There was no increase in the performance of DCRL-FSOA after two iterations on the validation set. The value of $t$ is set based on the distribution of output scores from DCRL-FSsev on the complete training set. Using the anomaly score in the 95% percentile resulted in the optimal performance on the validation set and therefore, this method for defining the threshold was implemented.
Margin, $m$.   The code implementation of this method automatically selects the margin value of $m$ based on the number of desired pseudo labels. The optimal performing number of pseudo labels on the validation set was $N=30$ for DCRL-FSOA and $N=3$ for DCRL-FSsev, resulting in margin values of 1.184 and 3.122 respectively. Keeping the margin static at 1.184 for the remainder of training resulted in 263 pseudo labelled anomalies in the second iteration of Stage 3 DCRL-FSOA.
Training Data Size.   Table 2 summarises the training data that was available at each stage and the hyper-parameters. Based on the analysis shown in section 8.3, the labelled training set size, $X_{train}$ was equal to 150 examples. $K=10$ ensembles were trained with $N=30$ for each ensemble as per the original FewSOME implementation. In the second/final iteration of Stage 3, DCRL-FSOA, there was substantially more data available for training due to the 263 pseudo labelled anomalies. To avoid a large class imbalance, the number of training instances for each ensemble was increased from $30$ to $150$ (all available training data).
Early Stopping.   For Stage 1, FewSOME’s early stopping method of halting training once the training loss began to plateau was employed. For all implementations in Stage 3, the training was stopped when the CD between the normal centre, $C_{norm}$ and anomalous centres $C_{OA}$ (when training DCRL-FSOA) or $C_{sev}$ (when training DCRL-FSsev) begins to plateau. Section 8.4 presents an analysis of the effects of this early stopping technique.
Table 2: The number of training examples available at each stage and hyper-parameters. Stage 3 is abbreviated to S3.
Stage # Labelled # Pseudo Architecture K m
Normal Labelled
Samples Anomalies
Stage 1 150 0 AlexNetl=5 10 $\times$
S3, DCRL-FSsev 150 3 VGG-16 10 3.122
S3iter1, DCRL-FSOA 150 30 VGG-16 10 1.184
S3iterfinal, DCRL-FSOA 150 263 VGG-16 10 1.184
@@ 5.1.2 Stage 3 Final Iteration
6 Evaluation
The methodology is assessed based on its ability to diagnose OA according to the KL scale (grades $\geq 2$) and the OARSI scale. The Area Under the Receiver Operating Characteristic Curve (AUC) is used to evaluate both tasks. The AUC is calculated by comparing the output of $S_{SSL}$ with the ground truth label for Stage 1. It is calculated by comparing the output of $S_{DCRL_{OA}}$, $S_{DRCRL_{sev}}$ and $S_{DRCRL_{comb}}$ with the ground truth label for Stage 3. The Spearman Rank Correlation Coefficient (SRC) between the model output and the KL grade is also reported. Finally, to ensure that the model can identify severe OA cases to a high degree of accuracy, the performance of the models on detecting cases with KL grades $>3$ is also calculated.
@@ 7 Results
Table 3 reports the AUC in % for the task of OA detection according to the KL and OARSI grading systems, denoted as $AUC_{KL}$ and $AUC_{O}$ respectively. The SRC between the model output and the KL grade is also reported as $SRC_{KL}$. The table also reports the performance of detecting cases with KL grades $>3$ as $AUC_{KL_{g>3}}$. The results were averaged across five seeds and one standard deviation is reported.
@@ 7.1 Stage 1
Each method had $150$ training instances available for training. The proposed SS-FewSOME significantly outperforms competing methods and a paired t-Test shows a statistically significant difference between the performance of the original FewSOME implementation and the proposed, increasing the $AUC_{KL}$ from $58.3$% to $72.9$%. The results also show that integrating the patches method into the model has resulted in a performance improvement, increasing $SRC_{KL}$ from 0.315 to 0.432.
@@ 7.2 Stage 3iter1
The continuous disease severity grading system, DCRL-FSOA improves the SRC from 0.432 in Stage 1 to 0.539 in Stage 3iter1. The trade-off between OA detection and severe OA detection accuracy can be evidently seen from the results of DCRL-FSOA (stage $3_{iter1}$) and DCRL-FSsev, with the former outperforming DCRL-FSsev in terms of OA detection but underperforming in terms of severe OA detection. However, the combined model DCRL-FScomb (stage $3_{iter1}$) combines both aspects of the models to accurately detect OA, severe OA and correlate with the KL grades with an SRC of 0.541. Figure 6 shows the distance of each data instance in the test set from $C_{norm}$ at stage 1 and stage 3iter1. Separation between classes can be evidently seen at stage 1 and furthermore at stage 3iter1. Figure 7 presents the most anomalous examples as output by DCRL-FScomb (stage $3_{iter1}$). The common symptoms of OA are apparent such as JSN and the presence of osteophytes.
@@ 7.3 Stage 3iterfinal
The final iteration achieved an SRC of 0.583, a promising result given the inter-observer reliability of human experts was reported to be as low as 0.51 in some cases [45]. The proposed methodology outperforms the competing method DDAD [12] by margins of up to 24%, which was trained on $150$ labelled healthy knee X-rays and 5,628 unlabelled X-rays. A paired t-Test showed this to be a statistically significant difference in performance. Additionally, a Two-Sample t-Test showed that the difference between the mean model output for each KL grade was statistically significant at the 5% significance level.
@@ 8 Further Analysis
8.1 CLIP Analysis
8.1.1 Compositional Statement Ablation
This analysis presents the metal detection performance of CLIP on the validation data for variations of compositional statements. The similarity score between each statement and the image in the validation set was calculated. The AUC was calculated between the output similarity scores and a label that indicates whether there was metal present in the image or not. The statement "there is a screw present in the image" resulted in the optimal performance with 99.3%.
Table 4: Comparison of the performance of CLIP on metal detection for variations of compositional statements.
Compositional Statement AUC (%)
"there is a screw present in the image" 99.3
"screw" 98.5
"there is a metal present in the image" 90.3
"metal" 94.0
"there is another object in the image that is not a knee X-ray" 2.0
8.1.2 CLIP Ablation
@@
DCRL-FSOA 3iter1 $\mathbf{78.6\pm{}1.4}$ $\mathbf{79.5\pm{}1.2}$ $\mathbf{95.2\pm{}1.5}$ $\mathbf{0.539\pm{}0.0}$
@@ 8.4 Analysis of Early Stopping
9 Discussion and Conclusion
This work has presented a three stage approach to developing a continuous disease severity grading system. The output anomaly scores of the proposed approach have a moderate to strong correlation with OA disease severity, achieving an SRC value of 0.58. Additionally, the proposed model can detect OA with an AUC of 0.81 according to both the KL and OARSI scales. The severe OA detection performance of the continuous grading system, DCRL-FSOA can be improved from 95.7% to 97.6% by combining it with the severe OA detector DCRL-FSsev, to form DCRL-FScomb. Although this causes a discontinuity in scores, the ranking of severity is preserved. The analysis showed that the proposed method outperforms existing methods by significant margins and can achieve human level performance whilst training on few labels.
The advantages of this approach over existing approaches fully supervised approaches are three fold. Firstly, the proposed method eliminates the requirement for substantial amounts of labelled data, reducing the barriers to clinical implementation. Secondly, this method is not trained to mimic a subjective grading system. Thirdly, diagnosing on a continuous grading system holds promise for substantial clinical impact by enabling the assessment of disease severity at any stage along a continuum, rather than predefined discrete categories. This method is particularly useful for capturing subtle variations and changes over time, providing a more detailed and dynamic understanding of the condition’s progression, ultimately resulting in improved patient care.
A limitation of this work is that its evaluation relies on comparisons with the KL grading system and the OARSI system, both of which are subjectively graded. The gold standard for evaluating the proposed system is assessment by clinicians. One potential method of evaluation would involve relative comparisons, where clinicians are presented with two knee X-rays and asked to determine which case has more severe OA. Their rankings could then be compared to the rankings generated by the proposed system. Additionally, clinicians could be asked to express their preference between the discrete KL grading system and the proposed continuous grading system for example cases. This would provide a high quality evaluation of the proposed system. An additional limitation is that the method requires the fine tuning of various hyper-parameters such as $m$ to labelled validation data.
Future work could explore additional datasets such as MOST [13] to further assess the generalisability of the approach. Additionally, the translation of such a system into clinical practice would require further work on explainability. Techniques such as Grad-CAM [38] could be employed to visualize the most salient regions of an image, providing insight into the system’s decision-making process. Additionally, a more comprehensive analysis would be needed to interpret what the values on the proposed scale represent in terms of OA features, such as the degree of JSN and the severity of osteophytes. Techniques such as Testing with Concept Activation Vectors (TCAV) [24] could be employed to quantify the importance of various visual biomarkers of OA in knee X-rays, thus improving the interpretability of the system.
@@ Acknowledgements
Appendix
Selection of $m$
This section presents the results on the validation set obtained from training on varying quantities of pseudo labelled training instances ($d$) at Stage 3 DCRL-FSOA. Table 7 shows in the AUC in % on the task of OA detection according to the KL grading system for values of $d$ ranging from 10 to 40. It can be observed from the table that the performance begins to plateau at $d=30$. As $d=30$ results in a class balance with the number of labelled training cases of healthy examples, $N=30$, $d$ was set to 30, resulting in margin $m$ value of 1.184. Table 8 shows the AUC in % for the task of severe OA detection at Stage 3 DCRL-FSsev. It can be observed that $d=3$ is the best performing implementation on the validation set, resulting in a value of $m$ of 3.122.
Table 7: AUC in % on the task of OA detection according to the KL grading system at Stage 3 DCRL-FSOA.
$d$ DCRL-FSOA
10 77.8
15 77.7
20 77.7
25 77.8
30 78.0
35 78.0
40 78.0
Table 8: AUC in % on the task of severe OA detection according to the KL grading system at Stage 3 DCRL-FSsev.
$d$ DCRL-FSsev
1 90.5
2 96.6
3 97.4
4 97.1
5 97.3